# GPU-Accelerated Gravitational Lensing &\&& Dynamical (GLaD) Modeling for Cosmology and Galaxies

### v1

@@ 2.4 Stellar dynamics
This equation describes stars embedded in a 3D gravitational field of the lens galaxy, with $\psi_{\rm D,int}$ being the deprojection of the 2D lensing potential $\psi_{\rm L,int}$ (up to a constant factor), ensuring phase-space density conservation. The phase-space density is not accessible for galaxies, and we can only measure the velocities $v$ along the LoS, and velocity dispersions $\sigma$ using the spectroscopy for distant galaxies $z>0.1$. To solve the Eq. 16, we reduce the number of the degree freedom by assuming an axisymmetric mass distribution ($\partial\psi_{\rm D,int}/\partial\phi=\partial f/\partial\phi=0$, with $\phi$ being the polar angle in the spherical coordinate system) and the spherically-aligned velocity ellipsoids. The choice of spherically-aligned velocity ellipsoids is due to the fact that lens galaxies are generally massive slow rotators. These galaxies exhibit a near-spherical mass distribution in their central regions, as opposed to a flat mass distribution characterized by cylindrically-aligned velocity ellipsoids. We multiply velocities along radial $v_{r}$, polar $v_{\theta}$ and azimuthal direction $v_{\phi}$ with Eq. 16 and integrate over all velocity space, obtaining two Jeans equations (e.g., Bacon et al. 1983, eqs. 1, 2)
@@
To derive the LoS velocities $\overline{v_{\rm LoS}^{2}}$ from Jeans equations (see Eqs. 17 and 18), it is essential to reconstruct the intrinsic luminosity and mass density of the lens galaxy in 3D. It is a common strategy to first apply multiple gaussian expansion (MGE Emsellem et al. 1994; Cappellari 2002) to the observed 2D surface brightness (SB) and mass convergence then deproject them later using inclination angle. The observed 2D SB of the lens galaxies, $I(x^{\prime},y^{\prime})$  [^6] Note that we present the general case here. In this paper, we perform 1D MGE fitting (see Sect. 3.2) to model the profile along the radial direction $R$ with a fixed axis ratio $q$, i.e., $R=\sqrt{x^{2}+y^{2}/q^{2}}$., is expressed through multiple Gaussians:
The deprojection process depends on the assumption of the galaxies’ shapes. For the commonly found elliptical galaxies with oblate shape, the deprojection requires:
$\cos^{2}{i}<q_{\rm min}^{\prime 2}$ (22)
where $r$ is the 3D radial distance to the galaxy centroid, $\theta$ is the polar angle (see definition in Cappellari 2020, Fig. 1), $\sigma_{j}=\sigma^{\prime}_{j}$ and $q_{j}=\frac{\sqrt{q^{\prime 2}_{j}-\cos{i}^{2}}}{\sin{i}}$ denote the dispersion and axis ratio of Gaussians after deprojection. The potential $\psi_{\rm D,int}$ in Eqs. 17 and 18 is derived by integrating the MGE of the 3D mass density $\rho_{\rm int}$. Following the approach used to infer the light tracer $\rho_{*}$, the 3D density profile $\rho_{\rm int}$ is obtained by deprojecting $\Sigma_{\rm int}$ (see Eq. 15). The surface mass density $\Sigma_{\rm int}$ is expressed as a sum of multiple Gaussian components,
Note we use index $i$ to denote the MGE components of the mass density, and $j$ for the luminosity components. The set of Gaussians describing the SB of lens galaxies (see Eqs. 21,  24) are not necessary identical to the MGEs of their mass densities. Therefore, $i\neq j$ meaning that $\sigma^{\prime}_{i}\neq\sigma^{\prime}_{j}$, $q^{\prime}_{i}\neq q^{\prime}_{j}$ and $M\neq N$ unless mass follows light. The deprojected $\rho_{\rm int}(r,\theta)$ is
The MGEs of $\rho_{*}(r,\theta)$ and $\rho_{\rm int}(r,\theta)$ are then substituted into the Jeans equations (17) and (18) to derive the intrinsic second velocity moments $\overline{v_{r}^{2}}$, $\overline{v_{\theta}^{2}}$, and $\overline{v_{\phi}^{2}}$. These moments correspond to the diagonal elements of the second velocity moment tensor, indicating a spherically aligned velocity ellipsoid, as all off-diagonal elements vanish.
@@ 3.1.1 Lensing modeling
The pixelated source $\bm{s}$ is reconstructed by maximising the posterior probability of $\bm{s}$, given the data
@@ 3.1.2 Dark matter profile $\kappa_{\rm enfw}$
We implement a dark matter profile following Oguri (2021) on the GPU, directly introducing ellipticity into the density mass profile $\kappa_{\rm eNFW}$, in contrast to the classical approach, which incorporates ellipticity in the potential. Since all lensing properties of $\kappa_{\rm eNFW}$ have analytical expressions, computing $\kappa_{\rm eNFW}$ and $\bm{\alpha}_{\rm eNFW}$ on a large grid of approximately $\mathcal{O}(10^{3})\times\mathcal{O}(10^{3})$ takes a negligible amount of time, approximately $10^{-5}\leavevmode\nobreak\ \text{sec}$ on a GPU. In contrast, performing the same computation on a CPU, following the approach of Golse & Kneib (2002), takes approximately 7 seconds. The detailed expressions for $\bm{\alpha}_{\rm eNFW}$ and $\psi_{\rm eNFW}$ are provided in Appendix A.
@@ 3.2 GPU acceleration in dynamical modeling
As discussed in Sect. 2.4, the MGE is commonly used in dynamical modeling as a prerequisite for JAM. Without accounting for the internal MSD, the surface brightness (SB) and mass density of the lens galaxies are sufficient for decomposition up to $3r_{\rm eff}$ in dynamic modeling. However, when considering the internal MSD, which represents a constant mass sheet added to the galaxy mass distribution, this additional mass can extend over a significantly larger region. To accurately account for the internal MSD, the mass profile must be decomposed over a larger area, approximately $\sim 50\arcsec$ for lens system RXJ1131 (Yıldırım et al. 2023; Shajib et al. 2023).
In Yıldırım et al. (2023), the authors applied the 2D MGE fitting method (Cappellari 2002) [^8] The adopted approach is the function mge_fit_sectors from the MgeFit package (https://pypi.org/project/mgefit/). to model the light and mass convergence map of the lens galaxy. In both cases, the maps are characterized by smooth profiles such as Sérsic, power-law, and NFW profiles, without any subtle angular structures. Since the maps primarily describe variations with radius, applying the 2D MGE fitting method is unnecessary in this case. The 2D MGE fitting method requires solving a non-linear least-squares minimization problem, which becomes computationally expensive when performed over a broad region extending $\sim 50\arcsec$ from the lens galaxy center. Moreover, producing the light and mass convergence maps in 2D across a wide area with $\mathcal{O}(10^{3})\times\mathcal{O}(10^{3})$ pixels is rather time-consuming, In total, it takes $\mathcal{O}(10)$ s per sampling step. The MGE 2D fit is primarily used to capture more detailed structures in galaxies from optical imaging directly, rather than relying on maps derived from profiles.
@@
We reimplement part of the jam.axi.proj function from the JamPy package [^9] https://pypi.org/project/jampy/ to compute $\overline{v_{\rm LoS}^{2}}$, the second velocity moment along the $z^{\prime}$-axis on the plane of the sky. The main computational bottleneck lies in solving the Jeans equations (Eqs. 17 and 18) to derive $\overline{v_{r}^{2}}$, $\overline{v_{\theta}^{2}}$, and $\overline{v_{\phi}^{2}}$ (see Sect. 5.1 in Cappellari (2020)). These computations involve numerical integrals, which is evaluated using adaptive quadrature methods in Shampine (2008). The integration region is initially divided into four subrectangles, and the integral in each subregion is computed using Gauss-Kronrod quadrature. If the estimated error in any subregion exceeds a predefined threshold, that subregion is further subdivided into four smaller subrectangles, and the process is repeated iteratively until the desired accuracy is achieved.
@@ 3.3 Joint modeling
In this section, we provide a detailed description of the joint modeling approach for time-delay cosmography. The input data $\bm{d}_{\rm LD}$ consist of both lensing and kinematic observations. The lensing data include the lens light, quasar image positions, the extended image of the host galaxy and the time delays between multiple observed images. The kinematic data comprise the spatially resolved kinematics map of the lens galaxy.
@@
We adopt parameterised mass profiles $\kappa_{\rm int}$ in the joint modeling. There are two mass classes
@@
In the first scenario, we model the baryonic component and dark matter of the lens galaxies separately. The baryonic component is represented by scaling the lens light profile $I_{\rm light}$, with a constant factor $\Upsilon_{\ast}$, while the dark matter is modeled using $\kappa_{\rm enfw}$ (see Eq. 62). $I_{\rm light}$ consists of two Chameleon profiles. The BH mass is included as a point mass $\kappa_{\rm BH}$. In the second scenario, we use an elliptical power-law (EPL) profile $\kappa_{\rm epl}$ to represent the total mass (see Appendix B). Because the EPL profile has a softening scale $r_{\rm scale}=0.01\arcsec$ that is set to a small value, the mass distribution diverges in the center, eliminating the need to add a separate point mass to represent the BH. In addition, we adopt an external shear to account for the tidal stretch from neighboring galaxies with external potential, expressed in polar coordinates $(R,\phi)$ as
@@
where $a_{0}$ is a normalisation parameter and $R^{2}=x^{2}+y^{2}$. In the region where $R\ll r_{\rm core}$, we obtain an approximate constant mass sheet
@@
In the region where $R\gg r_{\rm tr}$, we have $\lambda_{\rm int}\simeq 1$, indicating that the added mass sheet effectively vanishes at large scales. We emphasize that the values of $r_{\rm core}$ and $r_{\rm tr}$ are carefully selected based on extensive testing to represent the worst-case scenario. While the internal MSD remains unaffected from a lensing perspective, its impact on the kinematic data is significant enough to impose constraints on $\lambda_{\rm int}$. In addition, the dPIE profile, which has a well-defined truncation radius, declines more rapidly than the mass-sheet profile used in Blum et al. (2020). This makes it a more suitable choice, as it may help prevent negative densities in the outermost regions.
@@
Using the chosen mass density model, either a composite or power-law model, along with $I_{\rm light}$, we perform lensing and dynamical modeling simultaneously (see Fig. 2). Both the light and mass density profile of the lens galaxy must have the same position angle $\varphi_{\rm PA}$, to maintain the axisymmetric assumption. In our joint modeling, we fix this position angle to the mock input value. On the lensing side, we model the extended arc, lensing light, image positions, and time delays. For dynamical modeling, we decompose $I_{\rm light}$ and $\Sigma_{\rm int}$ into multiple Gaussian components. The MGE is carried out up to $50\arcsec$ from the lensing centroid, corresponding to approximately 200 kpc, ensuring that the mass density $\kappa_{\rm int}$, transformed by the internal mass sheet, remains physically meaningful at large distances. We focus on scenarios where the total mass density remains positive everywhere, ensuring physically valid predictions for $\bm{v_{\rm rms}^{\rm pre}}$, as negative densities would lead to unphysical results. To compute the predicted kinematic map, we incorporate the MGEs of $I_{\rm light}$ and $\Sigma_{\rm int}$ into the JAM modeling framework (see Sect. 2.4) to calculate $\bm{v_{\rm rms}^{\rm pre}}$. In practice, the light $I_{\rm light,IFU}$ near the spectral absorption lines in the IFU data should be provided to JAM to trace the stellar population responsible for these lines. In this paper, we work on the simulated kinematic data. However, we instead use the best-fit lens light model from the F814W filter in the infrared band. Since the lens galaxy in RXJ1131 is an early-type elliptical galaxy, the infrared band effectively characterizes the dominant stellar populations.
@@
where $\bm{d}_{\rm L}$ presents the lensing data, $\bm{d}_{\rm D}$ the kinematic data and $P(\bm{\eta}_{\rm LD})$ the prior for the lensing and dynamical parameters. The goodness-of-fit for a model is defined as
@@
We achieve a 20× speedup per sampling step using JAX on a single A100 GPU. Tab. 2 presents the runtime for each step using a composite mass model. Additionally, we include the runtime of the JAX code on a CPU for readers interested in evaluating the parallelization performance gains of JAX in different hardware. We note that the JAX is primarily optimized for GPU. On CPUs, its compilation overhead, lack of CPU-specific optimizations, and execution graph transformations can make it slower than NumPy.
@@
Table 2: Time comparison for one-step sampling in joint modeling using a composite mass model with a $64\times 64$ source grid (see Sect. 3.3 for the adopted profiles in the composite mass model). The computations were performed on a 2.10 GHz CPU and an NVIDIA A100 GPU. This table presents the runtime for the MGE 1D fit, both for a single mass or light profile (denoted as “1 profile”) and for the decomposition of all mass and light profiles in the modeling (denoted as “total”). The comparison of all MGE 1D fits is conducted using the same number of 21 Gaussians and the same number of log radii. The running time of the integral solver shows the calculation time for the second velocity moment on the diagonal of the tensor, which is the most time-consuming part for deriving $\bm{v_{\rm rms}^{\rm pre}}$. We adopt the same number of Guassins for testing, i.e. 42 Gaussins for lens light and 95 Gaussians for composite mass model. Note that JAX is primarily designed to maximize parallelization performance on GPUs. We present the running time of JAX code on a CPU to isolate the impact of GPU acceleration. In practice, the code is intended to run on GPUs.

### v2

@@ 2.4 Stellar dynamics
This equation describes stars embedded in a 3D gravitational field of the lens galaxy, with $\psi_{\rm D,int}$ being the deprojection of the 2D lensing potential $\psi_{\rm L,int}$ (up to a constant factor), ensuring phase-space density conservation. The phase-space density is not accessible for galaxies, and we can only measure the velocities $v$ along the LoS, and velocity dispersions $\sigma$ using the spectroscopy for distant galaxies $z>0.1$. To solve the Eq. 17, we reduce the number of the degree of freedom by assuming an axisymmetric mass distribution ($\partial\psi_{\rm D,int}/\partial\phi=\partial f/\partial\phi=0$, with $\phi$ being the azimuthal angle in the spherical coordinate system) and the spherically-aligned velocity ellipsoids. The choice of spherically-aligned velocity ellipsoids is due to the fact that lens galaxies are generally massive slow rotators. These galaxies exhibit a near-spherical mass distribution in their central regions, as opposed to a flat mass distribution characterized by cylindrically-aligned velocity ellipsoids. We multiply velocities along radial $v_{r}$, polar $v_{\theta}$ and azimuthal direction $v_{\phi}$ with Eq. 17 and integrate over all velocity space, obtaining two Jeans equations (e.g., Bacon et al. 1983, eqs. 1, 2)
@@
To derive the line-of-sight velocity moments $\overline{v_{\rm LoS}^{2}}$ from the Jeans equations (Eqs. 18 and 19), it is essential to reconstruct the intrinsic 3D luminosity and mass density distributions of the lens galaxy. It is a common strategy to first apply multiple Gaussian expansion (MGE; Emsellem et al. 1994; Cappellari 2002) to the observed 2D surface brightness (SB) and mass convergence, then deproject them later using the inclination angle (see Appendix. A). This yields the MGEs of the 3D luminosity density $\rho_{*}(r,\theta)$ and the 3D total mass density $\rho_{\rm int}(r,\theta)$. They are then substituted into the Jeans equations (18) and (19) to compute the intrinsic second velocity moments $\overline{v_{r}^{2}}$, $\overline{v_{\theta}^{2}}$, and $\overline{v_{\phi}^{2}}$. These moments correspond to the diagonal elements of the velocity dispersion tensor, assuming a spherically aligned velocity ellipsoid where all off-diagonal elements vanish.
@@ 3.1.1 Lensing modeling
The pixelated source $\bm{s}$ is reconstructed by maximizing the posterior probability of $\bm{s}$, given the data.
@@ 3.1.2 Dark matter profile $\kappa_{\rm enfw}$
We implement a dark matter profile following Oguri (2021) on the GPU, directly introducing ellipticity into the density mass profile $\kappa_{\rm eNFW}$, in contrast to the classical approach, which incorporates ellipticity in the potential. Since all lensing properties of $\kappa_{\rm eNFW}$ have analytical expressions, computing $\kappa_{\rm eNFW}$ and $\bm{\alpha}_{\rm eNFW}$ on a large grid of approximately $\mathcal{O}(10^{3})\times\mathcal{O}(10^{3})$ takes a negligible amount of time, approximately $10^{-5}\leavevmode\nobreak\ \text{sec}$ on a GPU. In contrast, performing the same computation on a CPU, following the approach of Golse & Kneib (2002), takes approximately 7 seconds. The detailed expressions for $\bm{\alpha}_{\rm eNFW}$ and $\psi_{\rm eNFW}$ are provided in Appendix B.
@@ 3.2 GPU acceleration in dynamical modeling
As discussed in Sect. 2.4, the MGE is commonly used in dynamical modeling as a prerequisite for JAM. Without accounting for the internal MSD, the SB and mass density of the lens galaxies are sufficient for decomposition up to $3r_{\rm eff}$ in dynamic modeling. However, when considering the internal MSD, which represents a constant mass sheet added to the galaxy mass distribution, this additional mass can extend over a significantly larger region. To accurately account for the internal MSD, the mass profile must be decomposed over a larger area, approximately $\sim 50\arcsec$ for lens system RXJ1131 (Yıldırım et al. 2023; Shajib et al. 2023).
In Yıldırım et al. (2023), the authors applied the 2D MGE fitting method (Cappellari 2002) [^6] The adopted approach is the function mge_fit_sectors from the MgeFit package (https://pypi.org/project/mgefit/). to model the light and mass convergence map of the lens galaxy. In both cases, the maps are characterized by smooth profiles such as Sérsic, power-law, and NFW profiles, without any subtle angular structures. Since the maps primarily describe variations with radius, applying the 2D MGE fitting method is unnecessary in this case. The 2D MGE fitting method requires solving a non-linear least-squares minimization problem, which becomes computationally expensive when performed over a broad region extending $\sim 50\arcsec$ from the lens galaxy center. Moreover, producing the light and mass convergence maps in 2D across a wide area with $\mathcal{O}(10^{3})\times\mathcal{O}(10^{3})$ pixels is rather time-consuming. In total, it takes $\mathcal{O}(10)$ s per sampling step. The MGE 2D fit is primarily used to capture more detailed structures in galaxies from optical imaging directly, rather than relying on maps derived from profiles.
@@
We reimplement part of the jam.axi.proj function from the JamPy package [^7] https://pypi.org/project/jampy/ to compute $\overline{v_{\rm LoS}^{2}}$, the second velocity moment along the $z^{\prime}$-axis on the plane of the sky. The main computational bottleneck lies in solving the Jeans equations (Eqs. 18 and 19) to derive $\overline{v_{r}^{2}}$, $\overline{v_{\theta}^{2}}$, and $\overline{v_{\phi}^{2}}$ (see Sect. 5.1 in Cappellari (2020)). These computations involve numerical integrals, which are evaluated using adaptive quadrature methods in Shampine (2008). The integration region is initially divided into four subrectangles, and the integral in each subregion is computed using Gauss-Kronrod quadrature. If the estimated error in any subregion exceeds a predefined threshold, that subregion is further subdivided into four smaller subrectangles, and the process is repeated iteratively until the desired accuracy is achieved.
@@ 3.3 Joint modeling
In this section, we provide a detailed description of the joint modeling approach for time-delay cosmography. The input data $\bm{d}_{\rm LD}$ consists of both lensing and kinematic observations. The lensing data include the lens light, quasar image positions, the extended image of the host galaxy and the time delays between multiple observed images. The kinematic data comprise the spatially resolved kinematics map of the lens galaxy.
@@
We adopt parameterised mass profiles $\kappa_{\rm int}$ in the joint modeling. There are two mass classes.
@@
In the first scenario, we model the baryonic component and dark matter of the lens galaxies separately. The baryonic component is represented by scaling the lens light profile $I_{\rm light}$, with a constant factor $\Upsilon_{\ast}$, while the dark matter is modeled using $\kappa_{\rm enfw}$ (see Eq. 63). $I_{\rm light}$ consists of two Chameleon profiles. The BH mass is included as a point mass $\kappa_{\rm BH}$. In the second scenario, we use an elliptical power-law (EPL) profile $\kappa_{\rm epl}$ to represent the total mass (see Appendix C). Because the EPL profile has a softening scale $r_{\rm scale}=0.01\arcsec$ that is set to a small value, the mass distribution diverges in the center, eliminating the need to add a separate point mass to represent the BH. In addition, we adopt an external shear to account for the tidal stretch from neighboring galaxies with external potential, expressed in polar coordinates $(R,\phi)$ as
@@
where $a_{0}$ is a normalisation parameter and $R^{2}=x^{2}+y^{2}$. In the region where $R\ll r_{\rm core}$, we obtain an approximately constant mass sheet
@@
In the region where $R\gg r_{\rm tr}$, we have $\lambda_{\rm int}\simeq 1$, indicating that the added mass sheet effectively vanishes at large scales. We emphasize that the values of $r_{\rm core}$ and $r_{\rm tr}$ are carefully selected based on extensive testing to represent the worst-case scenario. While the internal MSD remains unaffected from a lensing perspective, its impact on the kinematic data is significant enough to impose constraints on $\lambda_{\rm int}$. In addition, the dPIE profile, which features a well-defined truncation radius, declines more rapidly than the mass-sheet profile used in Blum et al. (2020). This makes it a more suitable choice, as it helps reduce the likelihood of having negative values for $\kappa_{\rm int}$ in the outermost regions. Negative mass convergence is unphysical and is therefore rejected by JAM.
Ideally, a mass profile for $\lambda_{\rm int}$ that characterizes the internal MSD for lens system RXJ1131 would remain constant out to $\sim 20\arcsec$, and then drop sharply to zero beyond that radius. This behavior would entirely prevent the emergence of negative $\kappa_{\rm int}$. However, such an “ideal” mass sheet lacks well-defined lensing properties and is difficult to fit using the MGE. In this work, we adopt a dPIE profile with a fixed $r_{\rm core}$ to represent the mass sheet. The relatively gradual decline of the dPIE profile in the outermost region and the fixed $r_{\rm core}$ can lead to negative $\kappa_{\rm int}$, resulting in an asymmetric probability distribution for $D_{\rm\Delta t,int}$ (see Sect. 5.2).
@@
Using the chosen mass density model, either a composite or power-law model, along with $I_{\rm light}$, we perform lensing and dynamical modeling simultaneously (see Fig. 2). Both the light and mass density profile of the lens galaxy must have the same position angle $\varphi_{\rm PA}$, to maintain the axisymmetric assumption. In our joint modeling, we fix this position angle to the mock input value. On the lensing side, we model the extended arc, lensing light, image positions, and time delays. For dynamical modeling, we decompose $I_{\rm light}$ and $\Sigma_{\rm int}$ into multiple Gaussian components. The MGE is carried out up to $50\arcsec$ from the lensing centroid, corresponding to approximately 200 kpc, ensuring that the mass density $\kappa_{\rm int}$, transformed by the internal mass sheet, remains physically meaningful at large distances. We restrict our analysis to models with strictly positive total mass density, as negative values would yield unphysical predictions for $\bm{v_{\rm rms}^{\rm pre}}$. The predicted kinematic map is then computed by feeding the MGEs of $I_{\rm light}$ and $\Sigma_{\rm int}$ into the JAM framework (see Sect. 2.4). In practice, the light $I_{\rm light,IFU}$ near the spectral absorption lines in the IFU data should be provided to JAM to trace the stellar population responsible for these lines. In this paper, we work on the simulated kinematic data. We instead use the best-fit lens light model from the F814W filter in the infrared band. Since the lens galaxy in RXJ1131 is an early-type elliptical galaxy, the infrared band effectively characterizes the dominant stellar populations.
@@
where $\bm{d}_{\rm L}$ presents the lensing data, $\bm{d}_{\rm D}$ the kinematic data, and $P(\bm{\eta}_{\rm LD})$ the prior for the lensing and dynamical parameters. The goodness-of-fit for a model is defined as
@@
We achieve a 20× speedup per sampling step using JAX on a single A100 GPU. Tab. 8 presents the runtime for each step using a composite mass model. Additionally, we include the runtime of the JAX code on a CPU for readers interested in evaluating the parallelization performance gains of JAX on different hardware. We note that JAX is primarily optimized for GPU. On CPUs, its compilation overhead, lack of CPU-specific optimizations, and execution graph transformations can make it slower than NumPy.
Table 2: Time comparison for one-step sampling in joint modeling
@@
888 Joint modeling is based on a composite mass model with a $64\times 64$ source grid (see Sect. 3.3 for the adopted profiles). The computations were performed on a 2.10 GHz CPU and an NVIDIA A100 GPU. This table presents the runtime for the MGE 1D fit, both for a single mass or light profile (denoted as “1 profile”) and for the decomposition of all mass and light profiles in the modeling (denoted as “total”). The comparison of all MGE 1D fits is conducted using the same number of 21 Gaussians and the same number of log radii. The running time of the integral solver shows the calculation time for the second velocity moment on the diagonal of the tensor, which is the most time-consuming part for deriving $\bm{v_{\rm rms}^{\rm pre}}$. We adopt the same number of Gaussians for testing, i.e. 42 Gaussians for lens light and 95 Gaussians for the composite mass model. Note that JAX is primarily designed to maximize parallelization performance on GPUs. We present the running time of JAX code on a CPU to isolate the impact of GPU acceleration. In practice, the code is intended to run on GPUs.